In how many different ways can 2 candidates be selected from a group of 25 applicants?

300

       
560

       
1140

       
2300
Answer #1

Solution

n C x = n! / x!(n - x)!

n! = n(n - 1)(n - 2) .... and so on .

n = 25 and x = 2

25 C 2 = 25! / 2!(25 - 2)! = 25! / 2!(23)! = 25 * 24 / 2 * 1 = 300

300 different ways